# Weight Space Representation Learning on Diverse NeRF Architectures

### v2

@@ Appendix B Additional Experimental Details
Encoder architecture. Our encoder is a slight re-adaptation of the graph meta-network used in Lim et al. [23] in their experiment called predicting accuracy for varying architectures. It has the same network hyperparameters: hidden dimension 128, 4 layers, a pre-encoder, 2 readout layers, and uses directed edges. However, unlike Lim et al., before average-pooling the edge features, it maps them to vectors of size 1024 through a single linear layer, resulting in a final embedding of size 1024 as in nf2vec [38].
@@
Decoder architecture. We use the nf2vec decoder in its original formulation [38]. The input coordinates $(x,y,z)$ are mapped to a vector of size 144 through a frequency encoding [26] and concatenated with the 1024-dimensional embedding produced by the encoder. The resulting vector of size 1168 becomes the input of the decoder, which consists of 4 hidden layers with dimension 1024 intertwined with ReLU activations, with a skip connection that maps the input vector to a vector of size 1024 via a single linear layer + ReLU and sums it to the output of the second hidden layer. Finally, the decoder outputs the four $(r,g,b,\sigma)$ predicted radiance field values.
@@
Table 8: NeRF classification (multi-architecture). Extended version of Table 3 with added $\mathcal{L}_{\text{C}}$ results. The encoder is trained on ALL; the classifier is trained on the datasets in column 2 and tested on those in columns 3–5.
Accuracy (%) $\uparrow$
Method Classifier Training Set MLP TRI HASH
nf2vec [38]
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) ALL 93.6 94.0 92.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 90.7 90.6 90.0
$\mathcal{L}_{\text{C}}$ (ours) 82.5 77.8 64.0
$\mathcal{L}_{\text{R}}$ (ours) MLP 93.8 25.3 19.3
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 91.5 58.6 56.6
$\mathcal{L}_{\text{C}}$ (ours) 83.2 60.9 38.6
$\mathcal{L}_{\text{R}}$ (ours) TRI 11.4 93.8 9.3
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 77.8 91.0 66.5
$\mathcal{L}_{\text{C}}$ (ours) 70.4 79.0 47.6
$\mathcal{L}_{\text{R}}$ (ours) HASH 13.8 35.7 92.7
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 54.1 35.5 90.6
$\mathcal{L}_{\text{C}}$ (ours) 65.4 63.7 64.6
Table 9: NeRF classification of unseen architectures (multi-architecture). Extended version of Table 4 with added $\mathcal{L}_{\text{C}}$ results. The encoder is trained on ALL; the classifier is trained on the datasets in column 2 and tested on those in columns 3–12, containing NeRF architectures unseen at training time.
@@
nf2vec [38]
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) ALL 91.3 87.4 93.2 88.3 24.8 69.4 91.9 91.2 88.3 24.5
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 85.9 83.8 87.0 84.1 72.1 30.8 89.2 87.4 86.8 27.8
$\mathcal{L}_{\text{C}}$ (ours) 67.3 47.7 63.5 43.3 60.4 4.6 51.9 43.5 55.9 27.7
$\mathcal{L}_{\text{R}}$ (ours) MLP 91.3 86.1 22.6 23.2 7.9 21.0 19.7 20.6 21.8 7.6
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 86.6 81.3 63.3 43.4 59.2 13.1 51.5 50.8 55.0 24.8
$\mathcal{L}_{\text{C}}$ (ours) 69.6 52.6 59.7 42.0 55.2 6.0 33.0 23.3 39.4 29.5
$\mathcal{L}_{\text{R}}$ (ours) TRI 11.9 10.6 92.1 84.8 32.1 43.6 9.2 8.8 15.0 5.2
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 60.7 58.0 86.9 83.2 63.9 30.2 59.8 61.5 66.3 28.5
$\mathcal{L}_{\text{C}}$ (ours) 54.6 37.5 62.8 44.9 57.3 5.8 36.2 27.9 43.1 28.4
$\mathcal{L}_{\text{R}}$ (ours) HASH 10.7 7.4 33.9 36.3 19.1 23.0 91.6 91.4 87.8 29.2
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 47.0 40.4 40.6 34.0 43.4 19.3 89.5 87.9 86.4 25.7
$\mathcal{L}_{\text{C}}$ (ours) 56.8 39.1 48.7 34.1 43.0 4.3 55.7 46.0 56.3 25.2
@@ Appendix C $\mathcal{L}_{\text{C}}$ Classification and Retrieval Results
NeRF classification. Tables 8 and 9 show classification results in the multi-architecture setting, including those produced by $\mathcal{L}_{\text{C}}$, which are omitted in the paper. Specifically, Table 8 shows classification results when the classifier $\mathcal{C}$ is tested on architectures seen during training, whereas Table 9 shows classification results when $\mathcal{C}$ is tested on architectures belonging to the families seen at training time but with different hyperparameters, which we refer to as unseen architectures in Section 4. In addition to the observations discussed in Section 4.2, Tables 8 and 9 show that $\mathcal{L}_{\text{C}}$ prevails over $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$ in very few cases; this behavior is consistent with the less noticeable separation between class-level clusters produced by $\mathcal{L}_{\text{C}}$ shown in Fig. 3 and discussed in Section 4.1. Notably, $\mathcal{L}_{\text{C}}$ in the best performing method when $\mathcal{C}$ is trained on HASH and tested on TRI in Table 8, where $\mathcal{L}_{\text{R}+\text{C}}$ struggles much more than in all other cases where $\mathcal{C}$ is tested on architectures that are not included in its training set (i.e. white cells).
NeRF retrieval. Table 10 shows retrieval results including those produced by $\mathcal{L}_{\text{C}}$, which are omitted in the paper. Since retrieval is an instance-level task, the contrastive loss allows $\mathcal{L}_{\text{C}}$ to perform better than $\mathcal{L}_{\text{R}}$ for every query/gallery combination, but not as well as $\mathcal{L}_{\text{R}+\text{C}}$, which always performs the best.
Table 10: Instance-level NeRF retrieval. Extended version of Table 6 with added $\mathcal{L}_{\text{C}}$ results. The encoder is trained on ALL. The reported query/gallery combinations belong to their corresponding test sets.
@@
nf2vec [38]
Cardace et al. [7]
$\mathcal{L}_{\text{R}}$ (ours) 1 0.0 1.8 0.4 4.5 1.7 1.1 0.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 5.3 30.6 14.8 33.3 13.2 13.5 9.5
$\mathcal{L}_{\text{C}}$ (ours) 2.9 16.4 6.2 15.9 4.3 5.4 4.0
$\mathcal{L}_{\text{R}}$ (ours) 5 0.0 5.3 1.8 13.3 6.1 3.8 1.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 17.7 59.4 38.0 61.2 33.4 32.1 25.1
$\mathcal{L}_{\text{C}}$ (ours) 10.9 40.6 18.0 38.9 14.1 16.8 14.2
$\mathcal{L}_{\text{R}}$ (ours) 10 0.0 8.2 3.1 19.2 9.2 6.1 2.7
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 27.2 72.7 51.3 72.6 45.0 42.8 36.3
$\mathcal{L}_{\text{C}}$ (ours) 18.0 53.8 26.9 53.6 22.0 26.3 21.9
@@ Appendix D Implementation and Hardware
Our framework implementation is built upon the codebases by Zama Ramirez et al. [38] (for decoder and training) and Lim et al. [23] (for the GMN architecture and graph conversion). NeRFs belonging to HASH and the unseen architectures were trained with the NerfAcc framework [22], as done in Zama Ramirez et al. [38] and Cardace et al. [7] for MLP and TRI. We plan to publicly release our code upon paper acceptance. All our experiments were performed on a single NVIDIA RTX A6000. Training either $\mathcal{L}_{\text{R}}$ or $\mathcal{L}_{\text{R}+\text{C}}$ took ${\sim}2$ weeks, whereas training $\mathcal{L}_{\text{C}}$ took ${\sim}4$ days.
@@ Appendix E Additional Qualitative Results
Figs. 6 and 7 show NeRF retrieval qualitative results produced by $\mathcal{L}_{\text{R}+\text{C}}$ for varying query/gallery combinations belonging to MLP, TRI, or HASH. It is worth noting that, although the low recall@$k$ in Table 10 might suggest an undesirable outcome, Figs. 6 and 7 reveal the similarity in color and/or appearance between the first 10 nearest neighbors and the query, which suggests that our framework was able to encode both relevant information about the appearance of the object represented by the NeRF and some invariance to the architecture used to parameterize it.
@@
Fig. 8 shows NeRF retrieval qualitative results produced by $\mathcal{L}_{\text{R}+\text{C}}$ when the query belongs to ObjaNeRF [3] and the gallery belongs to MLP, TRI, or HASH. As a reminder, ObjaNeRF’s NeRFs are vanilla MLPs with the same architecture as those in MLP (see Table 7). Since our framework was trained on ShapenetRender [37] NeRFs only, the purpose of this experiment is to shed some light on its ability to generalize to unseen datasets. The 10 nearest neighbors of the ObjaNeRF query reveal how remarkably well our method can retrieve NeRFs belonging to the same class as the query when the latter belongs to ShapenetRender classes (i.e. chair, lamp, and speaker), while mostly failing for the television query (whose closest ShapenetRender class would be display), although, interestingly, at least one object featuring a display with a pattern similar to that of the query is always found among the closest neighbors.
@@
Fig. 9 compares retrieval qualitative results of $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$ for one ObjaNeRF [2] query. As one may expect, the contrastive term in $\mathcal{L}_{\text{R}+\text{C}}$ allows retrieving objects belonging to the same class as the query even when the gallery NeRFs are parameterized by different architectures than the query (i.e. TRI and HASH), whereas $\mathcal{L}_{\text{R}}$ is able to produce reasonable results for all 10 nearest neighbors only when the gallery architecture is the same as the query (i.e. MLP), which, once again, demonstrates the importance of the contrastive objective when dealing with multiple architectures.
Figure 6: NeRF retrieval qualitative results part 1 ($\mathcal{L}_{\text{R}+\text{C}}$). Extended version of Fig. 4 (top). Query and gallery from MLP, TRI, or HASH.
Figure 7: NeRF retrieval qualitative results part 2 ($\mathcal{L}_{\text{R}+\text{C}}$). Extended version of Fig. 4 (top). Query and gallery from MLP, TRI, or HASH.
Figure 8: NeRF retrieval qualitative results ($\mathcal{L}_{\text{R}+\text{C}}$). Extended version of Fig. 4 (bottom). Query from ObjaNeRF [3], gallery from MLP, TRI, or HASH.

### v3

@@ Appendix B Additional Experimental Details
Encoder architecture. Our encoder is a slight re-adaptation of the graph meta-network used in Lim et al. (2024) in their experiment called predicting accuracy for varying architectures. It has the same network hyperparameters: hidden dimension 128, 4 layers, a pre-encoder, 2 readout layers, and uses directed edges. However, unlike Lim et al. (2024), before average-pooling the edge features, it maps them to vectors of size 1024 through a single linear layer, resulting in a final embedding of size 1024 as in nf2vec (Zama Ramirez et al., 2024).
@@
Decoder architecture. We use the nf2vec decoder in its original formulation (Zama Ramirez et al., 2024). The input coordinates $(x,y,z)$ are mapped to a vector of size 144 through a frequency encoding (Mildenhall et al., 2020) and concatenated with the 1024-dimensional embedding produced by the encoder. The resulting vector of size 1168 becomes the input of the decoder, which consists of 4 hidden layers with dimension 1024 intertwined with ReLU activations, with a skip connection that maps the input vector to a vector of size 1024 via a single linear layer + ReLU and sums it to the output of the second hidden layer. Finally, the decoder outputs the four $(r,g,b,\sigma)$ predicted radiance field values.
Training. Our encoder-decoder framework is trained end-to-end for 250 epochs with batch size 8, AdamW optimizer (Loshchilov & Hutter, 2019), one-cycle learning rate scheduler (Smith & Topin, 2019), maximum learning rate $1\text{\times}{10}^{-4}$ and weight decay $1\text{\times}{10}^{-2}$. These training hyperparameters are the same for $\mathcal{L}_{\text{R}}$, $\mathcal{L}_{\text{R}+\text{C}}$, and $\mathcal{L}_{\text{C}}$. In Eq.˜1, foreground and background pixels are weighted by 0.8 and 0.2, respectively, as done in Zama Ramirez et al. (2024). In Eq.˜3, $t$ and $b$ are initialized to $10$ and 10-10, respectively, as done in Zhai et al. (2023). In Eq.˜4, $\lambda$ is set to $2\text{\times}{10}^{-2}$, as we experimentally observed that this choice leads to $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{C}}$ values with the same order of magnitude.
Classification. In our classification experiments of Sec.˜4.2, the classifier is a concatenation of 3 (linear $\rightarrow$ batch-norm $\rightarrow$ ReLU $\rightarrow$ dropout) blocks, where the hidden dimensions of the linear layers are 1024, 512, and 256, and a final linear layer at the end computes the class logits. The classifier is trained via cross-entropy loss for 150 epochs with batch size 256, AdamW optimizer (Loshchilov & Hutter, 2019), one-cycle learning rate scheduler (Smith & Topin, 2019), maximum learning rate $1\text{\times}{10}^{-4}$, and weight decay $1\text{\times}{10}^{-2}$. These same network and training hyperparameters are used in the classification experiments of Zama Ramirez et al. (2024).
Table 15: NeRF classification (multi-architecture). Extended version of Tab.˜3 with added $\mathcal{L}_{\text{C}}$ results. The encoder is trained on ALL; the classifier is trained on the datasets in column 2 and tested on those in columns 3–5.
Accuracy (%) $\uparrow$
Method Classifier Training Set MLP TRI HASH
nf2vec (Zama Ramirez et al., 2024)
Cardace et al. (2024)
@@
[…]
Table 16: NeRF classification of unseen architectures (multi-architecture). Extended version of Tab.˜4 with added $\mathcal{L}_{\text{C}}$ results. The encoder is trained on ALL; the classifier is trained on the datasets in column 2 and tested on those in columns 3–12, containing NeRF architectures unseen at training time.
@@
nf2vec (Zama Ramirez et al., 2024)
Cardace et al. (2024)
[…]
@@ Appendix C $\mathcal{L}_{\text{C}}$ Classification and Retrieval Results
NeRF classification. Tabs.˜15 and 16 show classification results in the multi-architecture setting, including those produced by $\mathcal{L}_{\text{C}}$, which are omitted in Tabs.˜3 and 4. Specifically, Tab.˜15 shows classification results when the classifier $\mathcal{C}$ is tested on architectures seen during training, whereas Tab.˜16 shows classification results when $\mathcal{C}$ is tested on architectures belonging to the families seen at training time but with different hyperparameters, which we refer to as unseen architectures in Sec.˜4. In addition to the observations discussed in Sec.˜4.2, Tabs.˜15 and 16 show that $\mathcal{L}_{\text{C}}$ prevails over $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$ in very few cases; this behavior is consistent with the less noticeable separation between class-level clusters produced by $\mathcal{L}_{\text{C}}$ shown in Fig.˜3 and discussed in Sec.˜4.1. Notably, $\mathcal{L}_{\text{C}}$ in the best performing method when $\mathcal{C}$ is trained on HASH and tested on TRI in Tab.˜15, where $\mathcal{L}_{\text{R}+\text{C}}$ struggles much more than in all other cases where $\mathcal{C}$ is tested on architectures that are not included in its training set (i.e. white cells).
NeRF retrieval. Tabs.˜17 and 18 show retrieval results including those produced by $\mathcal{L}_{\text{C}}$, which are omitted in Tabs.˜7 and 7. The recall@$k$ of the Random baseline is computed as
$100\cdot\sum_{i=1}^{k}\frac{1}{|\text{gallery}|}=\frac{100\cdot k}{|\text{gallery}|}$ (5)
Since retrieval is an instance-level task, the contrastive loss enables $\mathcal{L}_{\text{C}}$ to outperform $\mathcal{L}_{\text{R}}$ for every query/gallery combination, albeit not as effectively as $\mathcal{L}_{\text{R}+\text{C}}$, which consistently yields the best results. In particular, when both the query and the gallery belong to ALL, $\mathcal{L}_{\text{R}}$ is unable to perform the task: the latent space organization shown in Fig.˜3 (left) and discussed in Sec.˜4.1 prevents $\mathcal{L}_{\text{R}}$ from being capable of recognizing the same object represented by NeRFs parameterized by different architectures. Furthermore, it is worth noting that the recall@$k$ in the ALL/ALL case is lower than in the other query/gallery combinations because the gallery is larger and the probability of missing the target instance is thus higher.
Table 17: NeRF retrieval (ShapenetRender). Extended version of Tab.˜7 with added $\mathcal{L}_{\text{C}}$ results. The encoder is trained on ALL. Query/gallery combinations belong to their respective test sets.
@@
nf2vec (Zama Ramirez et al., 2024)
Cardace et al. (2024)
Random 1 0.01 0.03 0.03 0.03 0.03 0.03 0.03
$\mathcal{L}_{\text{R}}$ (ours) 0.00 1.80 0.43 4.48 1.67 1.06 0.38
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 5.25 30.62 14.77 33.27 13.17 13.45 9.46
$\mathcal{L}_{\text{C}}$ (ours) 2.86 16.41 6.19 15.93 4.35 5.39 3.99
Random 5 0.04 0.13 0.13 0.13 0.13 0.13 0.13
$\mathcal{L}_{\text{R}}$ (ours) 0.00 5.28 1.82 13.30 6.12 3.84 1.37
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 17.72 59.37 37.98 61.24 33.40 32.09 25.13
$\mathcal{L}_{\text{C}}$ (ours) 10.91 40.61 18.03 38.89 14.11 16.84 14.21
Random 10 0.08 0.25 0.25 0.25 0.25 0.25 0.25
$\mathcal{L}_{\text{R}}$ (ours) 0.00 8.19 3.08 19.24 9.23 6.09 2.68
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 27.25 72.67 51.25 72.62 45.01 42.83 36.31
$\mathcal{L}_{\text{C}}$ (ours) 18.03 53.75 26.93 53.63 22.05 26.30 21.95
Table 18: NeRF retrieval (Objaverse generalization). Extended version of Tab.˜7 with added $\mathcal{L}_{\text{C}}$ results.. The encoder is trained on ALL. Query/gallery combinations belong to their respective test sets.
Recall@$k$ (%) $\uparrow$
nf2vec (Zama Ramirez et al., 2024)
Cardace et al. (2024)
Random 1 0.02 0.05 0.05 0.05 0.05 0.05 0.05
$\mathcal{L}_{\text{R}}$ (ours) 0.00 2.60 0.48 2.71 0.48 0.74 0.32
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 0.64 9.82 6.10 8.33 1.75 6.74 2.92
$\mathcal{L}_{\text{C}}$ (ours) 0.18 3.66 2.23 3.66 0.96 3.13 1.80
Random 5 0.09 0.27 0.27 0.27 0.27 0.27 0.27
$\mathcal{L}_{\text{R}}$ (ours) 0.00 6.37 1.11 7.27 1.91 2.92 0.69
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 2.69 23.67 17.94 20.70 5.73 19.59 8.49
$\mathcal{L}_{\text{C}}$ (ours) 1.65 13.22 8.39 10.08 3.18 10.46 5.47
Random 10 0.18 0.53 0.53 0.53 0.53 0.53 0.53
$\mathcal{L}_{\text{R}}$ (ours) 0.00 9.66 2.39 10.99 3.24 4.88 1.80
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 4.87 33.65 26.38 27.55 9.50 28.13 12.74
$\mathcal{L}_{\text{C}}$ (ours) 3.01 19.53 13.48 15.29 5.57 15.45 8.39
@@ Appendix D Implementation and Hardware
Our framework implementation is built upon the codebases by Zama Ramirez et al. (2024) (for decoder and training) and Lim et al. (2024) (for the GMN architecture and graph conversion). NeRFs belonging to HASH, $\texttt{TRI}^{\texttt{OB}}$, $\texttt{HASH}^{\texttt{OB}}$, and the unseen architectures were trained with the NerfAcc framework (Li et al., 2022), as done in Zama Ramirez et al. (2024) and Cardace et al. (2024) for MLP and TRI. We plan to publicly release our code and datasets upon paper acceptance. All our experiments were performed on a single NVIDIA RTX A6000. Training either $\mathcal{L}_{\text{R}}$ or $\mathcal{L}_{\text{R}+\text{C}}$ took ${\sim}2$ weeks, whereas training $\mathcal{L}_{\text{C}}$ took ${\sim}4$ days.
@@ Appendix E Additional Qualitative Results
Figs.˜6 and 7 show NeRF retrieval qualitative results produced by $\mathcal{L}_{\text{R}+\text{C}}$ for varying query/gallery combinations belonging to MLP, TRI, or HASH. The similarity in color and/or appearance between the first 10 nearest neighbors and the query suggests that our framework encodes both relevant information about the appearance of the object represented by the NeRF and some invariance to the architecture used to parameterize it.
@@
Fig.˜8 shows NeRF retrieval qualitative results produced by $\mathcal{L}_{\text{R}+\text{C}}$ when the query belongs to $\texttt{MLP}^{\texttt{OB}}$ (Amaduzzi et al., 2025) and the gallery belongs to MLP, TRI, or HASH. As a reminder, $\texttt{MLP}^{\texttt{OB}}$’s NeRFs are vanilla MLPs with the same architecture as those in MLP (see Tab.˜14). Since our framework was trained on ShapenetRender (Xu et al., 2019) NeRFs only, the purpose of this experiment is to shed some light on its ability to generalize to unseen datasets. The 10 nearest neighbors of the $\texttt{MLP}^{\texttt{OB}}$ query reveal how remarkably well our method can retrieve NeRFs belonging to the same class as the query when the latter belongs to ShapenetRender classes (i.e. chair, lamp, and speaker), while mostly failing for the television query (whose closest ShapenetRender class would be display), although, interestingly, at least one object featuring a display with a pattern similar to that of the query is always found among the closest neighbors.
@@
Fig.˜9 compares retrieval qualitative results of $\mathcal{L}_{\text{R}}$ and $\mathcal{L}_{\text{R}+\text{C}}$ for one $\texttt{MLP}^{\texttt{OB}}$ (Amaduzzi et al., 2025) query. As one may expect, the contrastive term in $\mathcal{L}_{\text{R}+\text{C}}$ allows retrieving objects belonging to the same class as the query even when the gallery NeRFs are parameterized by different architectures than the query (i.e. TRI and HASH), whereas $\mathcal{L}_{\text{R}}$ produces reasonable results for all 10 nearest neighbors only when the gallery architecture is the same as the query (i.e. MLP), which, once again, demonstrates the importance of the contrastive objective when dealing with multiple architectures.
Appendix F Additional NeRF Captioning and Q&A Results
Tabs.˜24, 24, 24, 24, 24 and 24 show extended results for the language tasks described in Sec.˜4.4 by adding BLEU-1 (Papineni et al., 2002), ROUGE-L (Chin-Yew, 2004), and METEOR (Banerjee & Lavie, 2005) metrics, which are traditional handcrafted similarity measures based on n-gram statistics. As discussed in Sec.˜4.4, S-BERT (Reimers & Gurevych, 2019) and SimCSE (Gao et al., 2021) are instead based on pre-trained networks, and generally considered more effective than n-gram-based metrics at measuring the quality of the generated text (Amaduzzi et al., 2024).
Table 19: NeRF brief captioning (multi-architecture). Extended version of Tab.˜13.
Method Test Dataset S-BERT $\uparrow$ SimCSE $\uparrow$ BLEU-1 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
LLaNA (Amaduzzi et al., 2024)
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) MLP 67.1 68.9 32.7 35.1 36.1
TRI 66.4 67.8 32.5 34.4 35.9
HASH 66.3 67.9 32.4 34.5 35.6
Table 20: NeRF detailed captioning (multi-architecture). Extended version of Tab.˜13.
Method Test Dataset S-BERT $\uparrow$ SimCSE $\uparrow$ BLEU-1 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
LLaNA (Amaduzzi et al., 2024)
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) MLP 73.3 75.1 19.7 32.1 20.1
TRI 72.2 73.8 19.5 31.4 19.7
HASH 72.4 74.0 19.6 31.7 19.9
Table 21: NeRF single-round Q&A (multi-architecture). Extended version of Tab.˜13.
Method Test Dataset S-BERT $\uparrow$ SimCSE $\uparrow$ BLEU-1 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
LLaNA (Amaduzzi et al., 2024)
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) MLP 80.9 81.4 45.6 52.6 49.5
TRI 80.7 81.3 45.7 52.9 49.8
HASH 80.6 81.3 45.3 52.4 49.3
Table 22: NeRF brief captioning (single-architecture). Extended version of Tab.˜13.
Method S-BERT $\uparrow$ SimCSE $\uparrow$ BLEU-1 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
LLaNA (Amaduzzi et al., 2024) 68.6 70.5 20.6 28.3 31.8
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 68.3 70.2 33.1 35.3 36.6
Table 23: NeRF detailed captioning (single-architecture). Extended version of Tab.˜13.
Method S-BERT $\uparrow$ SimCSE $\uparrow$ BLEU-1 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
LLaNA (Amaduzzi et al., 2024) 77.4 79.8 41.3 36.2 32.4
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 74.1 76.0 20.1 32.5 20.6
Table 24: NeRF single-round Q&A (single-architecture). Extended version of Tab.˜13.
Method S-BERT $\uparrow$ SimCSE $\uparrow$ BLEU-1 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
LLaNA (Amaduzzi et al., 2024) 81.0 81.6 46.2 53.2 50.2
$\mathcal{L}_{\text{R}+\text{C}}$ (ours) 81.0 81.6 45.4 52.5 49.4
Appendix G LLM Usage
LLMs were used in this paper as a tool to aid and polish writing. No significant role in research ideation, experimentation, or writing was delegated to LLMs.